What is the percentage by mass of the solution that contains 10.0 grams of Phosphorus pentachloride that has been dissolved in 750 mL of water?


1
Expert's answer
2022-02-19T06:51:08-0500

Mass of water =750ml×11g/ml=750g

Total mass = 750+10=780g

Percentage of PCl5 =(10÷780)×100

=1.282%

Percentage of water=98.718%